A solid conducting sphere of radius 20.00cm has a charge of 2.00ųC. If the potential is zero at infinity, find the value of the potential at the following distances from the center of the sphere:



a. 5.00cm b. 10.00cm c. 15.00cm

1
Expert's answer
2022-04-29T12:30:22-0400

The potential of electric field inside conductor is the same. Hence


"V_a=V_b=V_c=k\frac{q}{r}"

"=9*10^9*\frac{2.00*10^{-6}}{0.20}=90000\:\rm V"